### 14 Dec 2023 v1

# Google Tag Manager: Hidden Data Leaks and its Potential Violations under EU Data Protection Law

Abstract.
Tag Management Systems were developed in order to support website publishers in installing multiple third-party JavaScript scripts (Tags) on their websites. In 2012, Google developed its own TMS called “Google Tag Manager” (GTM) that is currently present on 28 million live websites. In 2020, a new “Server-side” GTM was introduced, allowing publishers to include Tags directly on the server. However, neither version of GTM has yet been thoroughly evaluated by the academic research community.
In this work, we study, for the first time, the two versions of the Google Tag Management (GTM) architectures: Client- and Server-side GTM. By analyzing these systems with 78 Client-side Tags, 8 Server-side Tags and two Consent Management Platforms (CMPs) from the inside, we discover multiple hidden data leaks, Tags bypassing GTM permission system to inject scripts, and consent enabled by default. With a legal expert, we perform an in-depth legal analysis of GTM and its actors to identify potential legal violations and their liabilities. We provide recommendations and propose numerous improvements for GTM to facilitate legal compliance.
online tracking, server-side tracking, privacy, consent, GDPR compliance, website publishers, data controller, potential legal violation
journalyear: YYYYjournalvolume: YYYYjournalnumber: Xdoi: XXXXXXX.XXXXXXX
1. Introduction
Today’s modern websites continuously collect their visitors’ data for various purposes, such as targeted advertising, and rely on third parties for such collection. Over the last decade, researchers have demonstrated that third-parties collect users’ data with the help of third-party JavaScript scripts (Nikiforakis et al., 2012). These scripts, invisible to the users, are silently executed in the webpage’s background and are often called “tags” by the web marketing industry (Google, ndo). Initially, to install a tag on the webpage, the website publisher only needed to copy and paste an external JavaScript library reference to the webpage’s source code. However, as publishers were installing more and more tags [^1] In 2016, Engelhard and Narayanan (Englehardt and Narayanan, 2016a) measured that on average, a website contains 18 third-parties., manual tag management became challenging and, as a result, “Tag Management Systems” (TMS) were developed by the industry. TMS allow publishers to install and configure tags in a centralized manner without tinkering with the website source code. Once installed on a website, TMS takes care of handling the installation, configuration and execution of third-party tags.
In 2012, Google developed its own TMS called “Google Tag Manager” (GTM). It is currently the most installed TMS on the market, currently present on 28 million live websites according to BuiltWith (BuiltWith® Pty Ltd, nd). GTM is a free service, offering a graphical interface and supporting a seamless inclusion of major marketing and analytic third-party scripts. GTM also benefits from a community of contributors, creating tags for services that are not officially supported by Google. Tags rely on the ability of browsers to communicate directly with third-party domains to download scripts, set cookies and send users’ data. Figure 0(a) shows the first GTM architecture proposed by Google that we call “Client-side GTM” since it loads all the tags inside the user’s browser. In the recent years, this architecture became crippled by measures taken by browser vendors. Many popular Web browsers, such as Safari, Firefox and Brave, already actively block third-party tracking scripts and cookies to defend users against Web tracking (John Wilander, 2019; Mozilla Corporation, nd; Brave Software, Inc, nd). Moreover, Google’s own plan of phasing out third-party cookies in Chrome (Justin Schuh, 2020) will render a lot of tags that rely on third-party cookies ineffective.
As a result, an alternative version of Google Tag Manager was created in 2020, officially called “GTM Server Container” (Google, ndn). This new architecture, shown in Figure 0(b), loads and executes tags in a remote server, making it look like no third parties are present on the website. When the user’s data is collected with the GTM Web Container, there is only one data flow exiting the browser, and therefore, the final destination of data is impossible to trace for researchers and auditors that analyze outgoing browser requests. This architecture, called “Server-side GTM”, bypasses any browser restrictions and security measures, such as CSP (mozilla.org contributors, nda) to control and secure third-party scripts and allows tags, that are now run on the server, to invisibly share users’ data to other third parties.
(a) Client-side GTM
(b) Server-side GTM
Figure 1. Google Tag Management (GTM) architecture. Red arrows represent the flows of users’ data to third parties.
In this work, we study two versions of the Google Tag Management (GTM) architectures – Client- and Server-side GTM – from inside. Differently from other approaches to measure Web tracking (Roesner et al., 2012; Lerner et al., 2016; Fouad et al., 2020a, 2022), we deploy the idea of Toth et al. (Toth et al., 2022) to analyze the system by installing it on an empty website that we fully control. We then perform two separate studies: Study I that focuses on the internal functioning of GTM and Study II that analyzes legal implications of Study I. In Study I (§3), our first contribution consists in the methodology we built to analyze the internals of Client- and Server-side GTM, by installing 78 Tags one-by-one and two popular consent banner solutions. This allows us to capture script injections, hidden data flows, study GTM permission system and its integrated “Consent Mode”. We discovered 6 technical findings, including the fact that “Google Tag” aims at ensuring communication between Client- and Server-side, collects multiple types of users’ data without their consent; “Pinterest Tag” collects a significant amount of users’ data without disclosing it to the Publisher; 11 out of 78 official Client-side tags inject a third-party script into the DOM (mozilla.org contributors, ndb) bypassing the GTM permission system; and GTM “Consent Mode” enables some of the consent purposes by default, even before the user has interacted with the consent banner.
Study II in §4 is written together with a legal expert and co-author of this paper. Therein we explain the EU Data Protection legal background relevant to the GTM ecosystem. We then provide an in-depth legal analysis where we identify the legal role of each actor, detect 8 potential violations of the General Data Protection Regulation (GDPR) (GDPR, 2016) and ePrivacy Directive (ePD-09, 2009a), and thereby we identify each actor’s liability. Finally, we make recommendations to further improve GTM to facilitate legal compliance across all involved actors. Finally, we discuss other potential legal implications of this GTM ecosystem (§5) and conclude the paper.
2. Related Works
GTM has become a popular tool for managing tags on websites in recent years. However, to the best of our knowledge, no scientific study has analyzed tag managers or the GTM Framework in particular, neither its client- nor its server-side tagging version from a privacy viewpoint. In this section we present the related works in the closest areas: online tracking detection and measurement, legal requirements on consent for online tracking, and non-academic literature, such as blog posts, regarding GTM.
Online tracking measurement and prevalence of Google Multiple works have detected and measured web tracking practices over the last decade, identifying Google’s massive prevalence in the tracking ecosystem. In 2009, Krishnamurthy and Wills (Krishnamurthy and Wills, 2009) conducted a longitudinal study of the collection and aggregation of personal data of end users by third parties – they highlight a small number of actors in the tracking ecosystem, and the rapid growth of Google, with requests to Google-owned servers being observed on 60% of the domains. In 2015, by analyzing third-party HTTP requests on the Alexa top 1 million sites, Libert (Libert, 2015) concluded that Google could track users on up to 80% of sites. In 2016, Englehardt and Narayanan (Englehardt and Narayanan, 2016b) confirmed the extent of this tracking by conducting an automated analysis of stateful and stateless techniques on 1 million sites, showing that Google owned the five most detected third-party domains. In 2020, Fouad et al. (Fouad et al., 2020a) confirmed again the prevalence of Google – they found the presence of tracking by Google domains on more than 85% websites.
Tag management and GTM No academic research appears to have studied tag management systems to date. However, several people close to AdTech circles share their experience and publish screenshots and other analyses of these tools on the web, either to help publishers in their deployment or to point out legal or technical issues. The IT and digital marketing expert Julius Fedorovicius publishes courses and ebooks on GTM on his website Analytics Mania (Julius Fedorovicius, nd). He has published a how-to on configuring server-side GTM (Julius Fedorovicius, 2023), summarizing the essential benefits and problems of the technology. The analytics developer Simo Ahava tests and comments on new functionalities in SEO and digital marketing tools. He maintains an extensive documentation on GTM (Simo Ahava, nd), and in a recent article (Simo Ahava, 2022), discusses GTM’s limitations and problems regarding transparency and user control. The blogger Pixel de tracking (PixeldeTracking-website, nd) explores surveillance issues on the web. They have published several articles on Web Tracking Technologies, including GA4 and server-side GTM, focusing on the implications on the impact on content blockers (de Tracking, 2020). A number of books have been published on the subject of GTM, attempting to guide website publishers in their use of the tool (Weber, 2015; Chardonneau, 2015). They cover registration, configuration, tagging and integration with popular CMS such as WordPress.
Summary Most of the previous works either focus on Web Tracking Technologies in general, or just describe the functionalities of GTM or server-side tag management. The recent content published by industry experts identify risks for privacy. However, no work so far seems to have evaluated if GTM can be leveraged to deploy offensive configurations at scale, nor seems to have assessed the complexity of consent management in GTM for publishers in the GDPR context. It is these gaps that our work aims to fill.
3. Study I: the internals of GTM
In Study I, we experimented with both the Client- and Server-side GTM to identify their components and how they work in a typical GTM installation. Our goal was to evaluate what and how user data is collected and which actors access it. We examined GTM’s “Consent Mode”, a toolset to render GTM compliant, in order to assess its effectiveness in managing consent.
3.1. Methodology
Tag name (ranking (Binns et al., 2018)) Data Collector Goals of the service/tag
Google Tag (1)
Google
provide statistics on users and their behavior on websites
Pinterest Tag (66)
Pinterest
measure advertisement campaigns effectiveness; become a Pinterest Verified merchant
Hotjar Tracking Code (105)
Hotjar
provide videos of user interactions with websites to detect bugs and improve the website
(a) Client-side GTM tags.
Tag name (ranking (Binns et al., 2018)) Data Collector Goals of the service/tag
Google Analytics: GA4 (1)
Google
provide statistics on users and their behavior on websites
Conversions API Tag (3)
Facebook
measure advertisement campaigns effectiveness; provide statistics about website usage
Mixpanel (32)
Mixpanel
gather data on user interactions with the website to gain insights into their behavior
(b) Server-side GTM tags.
Table 1. Tags selected for further study in Client- and Server-side GTM.
To conduct experiments and set up the GTM infrastructure, we bought a domain – we call it example.com here – and created a public website containing one basic webpage with a paragraph of text and an HTML login form. We have included a login form since Senol et al. (Senol et al., 2022) have recently found that user input is often leaked from the forms, so we decided to test whether Tags may be responsible for such leakage. The website and the Server-side GTM infrastructure were hosted on a virtual machine we rented on the Microsoft Azure cloud computing platform located in a data center in the EU. We conducted this study between September and November 2023 on the Chromium browser (chromium-browser, nd) in version 111, from the Flathub repository (chromium-flatpak, nd), using the default settings and installed on a GNU/Linux operating system (kernel version 6.1.x-lts). We used the “profiles” functionality of the browser to start every experiment in a fresh environment, devoid from cookies, local storage and other technologies than maintain a state. The browser, visiting the website, was run on a computer connected to the Internet through an institutional network in the EU. To create Client- and Server-side GTM installations, we created a new Google account, logged into it and followed the suggested steps in the official GTM documentation (Google, ndl). We explain the installation process in detail below. Note that the terms we use can slightly differ from that of the official GTM documentation for clarity reasons.
3.1.1. Client-side GTM
Selecting Tags. GTM supports 78 official Tags that are directly accessible through its interface to the website publisher. Nevertheless, a community of developers created Tags for services that are not officially supported by Google, and such tags are available through the GTM-integrated template gallery (Google, ndm). To understand what type of user’s personal data Tags can collect, and to monitor how it flows through the Client-side GTM installation, we decided to select and test 3 Tags. To select Tags, we decided to pick representative Tags of three typical functionalities of third parties: analytics/statistics, advertisement effectiveness, and assessment of user interaction. Our selection procedure focuses on the list of official Tags and was informed by the popularity of the company that receives the data collected by the Tag (we call such company Data Collector, see §4.2.3). For the popularity criteria, we followed the ranking of popularity of third parties according to Binns et al. (Binns et al., 2018), who measured the inclusion of third parties on 100k websites. The three selected Tags are shown in Table 0(a) [^2] For clarity, we will refer to the “Hotjar Tracking Code” tag as the “Hotjar Tag”.. For the selected Tags to function properly, we registered on the Data Collector’s website and configured the Tag according to its instructions.
Installing GTM on our Website. When the publisher has chosen and configured Tags to be included on its website, GTM generates a “Web Container” in the form of an external JavaScript library called gtm.js  (Google, ndf), which contains all the selected tags. We did it and copy-pasted the small script provided by GTM in our website source page, which prompts the browser to fetch the gtm.js script.
(a) Client-side GTM.
(b) Server-side GTM.
Figure 2. Pipelines of the analysis of collected data and tag behavior on Client- and Server-side GTM.
Capturing Script injection, Data Flows and Availability of Collected Data. We manually installed all the officially available tags individually in our Web Container, opened our website in a new browser profile and analyzed the traffic using the Chromium debug tools. For each tag we identified whether additional scripts were downloaded and for the 3 tags selected in Table 0(a), we inspected the GET parameters, POST bodies and data exchanged through the WebSocket protocol of outgoing requests to identify the data collected. In order to identify personal data collected, we investigate firstly by using an empirical method: by looking at the key/value pairs, in GET parameters or JSON data, in search for obvious key names like screen_resolution or obvious key/value pairs such as sr=1920x1080 for a FullHD screen; and secondly by using the available documentation for each Tag (Google, ndi; Pinterest, nda; Hotjar, nd). Finally, we logged on the Data Collector’s website, searched for the personal data previously identified and compared it with data identified in the outgoing requests. For the Google Tag, in Google Analytics, we visited the “Report” page; for Pinterest, in the “Conversion” section, we clicked on the “Event history” page; for Hotjar we went to the “Recordings” page to see the individual recorded user interactions. Figure 1(a) summarizes our analysis method.
Tag Permission System. For non-official tags, GTM shows the permissions of the tag in a popup, to review them before confirming the installation. By analyzing the network communications with our browser while being in the Web Container configuration interface, in the template gallery section, we found a JSON file that contains the tag information displayed in the interface. This file contains information such as the tag’s author, a description of the tag, the link to the source code of the tag and the list of permissions. The names of permissions and configurations (e.g., for the set_cookies permission, the configuration would be the cookie name) corresponds to the official documentation (Google, nde).
For official tags however, the GTM configuration interface does not show any permission and no official documentation mentions any permission system either. However, by analyzing network communications while browsing the official tags section, we found a similar JSON file containing the list of permissions in the same format. From this file we extracted the list of tags having the inject_script permission to compare it with the actual behavior of the tags regarding script injection. We conclude that official tags have a permission system similar to the non-official tags.
Consent Configuration. Next, assisted by a legal scholar coauthor of this paper, we integrated consent management on our website in such a way that it is compliant with the EU legal framework. We used the GTM “Consent Mode” feature, added in 2020 in Web Containers, and still marked as a beta feature in the GTM configuration interface during the writing of this paper. This consent system is based on “consent variables”, and GTM proposes: ad_storage , analytics_storage , functionality_storage , personalization_storage and security_storage (Table 3). They take two values, “granted” or “denied”, according to the documentation. We found that variables can also be in “undefined” state, with serious consequences (see Technical finding 4). In this system, the Consent Management Platforms (CMPs), which provides consent banners (Matte et al., 2020), are integrated in the Web Container as Tags. CMPs communicate the user’s consent choices made in the consent banner, expressed through these consent variables, to the Web Container (Google, ndc). Then, in the GTM interface, we, as Publisher, need to associate none, one or more consent variables to each Tag. According to documentation (Google, ndh), before executing a Tag, GTM checks the associated consent variables and proceeds only if all the consent variables are “granted”; otherwise GTM prevents the tag from running. Some tags support a feature called “built-in consent” (Google, ndd) which allows them, when being executed, to check the values of the consent variables by themselves and adapt their behavior depending on users’ choices. If trusted, the Publisher does not associate any consent variable with such tags that are always executed. We tested the two CMPs marked as compatible with Consent Mode and put forward in the GTM interface, namely Consentmanager (Consentmanager, nda) and Cookiebot (Cookiebot, 2020). They both include a scanner to detect cookies and/or tracking technologies and classify the trackers and associated purposes (we further analyze scanners in section 4.2.1). For each CMP, we created an account on their website, let the CMP scanner scan our website, and installed the CMP tag in the GTM configuration interface. For Consentmanager, according to the documentation, we enabled “Send Google Consent Mode” in the settings of the CMP’s website. We then used the results of the CMP scanner to configure the consent settings of each tag. To study the behavior of the Web Container, the CMP and the Tags, we visited our website several times, using a new browser profile and a different consent option in the consent banner each time: “accept all”, “decline all”, “only analytics” and “only advertisement”. After making our choice, we stayed on the page for 20 seconds, reloaded it and used it for another 20 seconds. We used both the debugger provided by GTM to know when tags run, and the browser’s debugger to capture outgoing traffic (figure 1(a)). Finally, we compared collected data in each case with data collected without consent configuration.
3.1.2. Server-Side GTM
Selecting tags. Similarly to Client-side GTM, we selected three tags with the same representative functionalities of third parties: analytics/statistics, advertisement effectiveness, and assessment of user interaction. Here, GTM officially supports only eight tags, all related to Google services, from which we selected Google Analytics. Then, from the template gallery where developers propose additional, non-officially supported tags, and for the remaining two functionalities we are interested in, we chose the most popular ones, based on the ranking of the Data Collector as in Binns et al. (Binns et al., 2018), namely Facebook Conversion API and Mixpanel (Table 0(b)). We also observed that in the template gallery, the company that provides the tag (we call it Tag Provider, see §4.2.4) and the Data Collector are not necessarily the same.
GTM Server Container Installation. In the GTM interface for Publishers, we created a “GTM Server Container”, which is a Node.js program running on a server, that contains in particular the selected “Server Tags”. Two options are proposed: “Automatic provisioning”, that automatically installs the Server Container on a Google Server, and “Manual installation”, for which the Publisher provides the server. We chose the manual installation to have total control over the server and be able to analyze the Server Container incoming and outgoing information flows. We used the provided Docker image and configured the Server Container to be accessible using HTTPS. Finally, we created A and AAAA DNS records for the gtm.example.com subdomain pointing to our server and avoided using CNAME records as recommended. We provide the configuration files of our setup in the artifacts (§3.1.3).
Connection Between the browser and the Server Container. To connect the browser and the Server Container to collect the end-user data, we first created a new GTM Web Container (figure 1(b)). Following the GTM documentation, we installed the “Google Tag” in the Web Container, specifying the https://gtm.example.com URL of our Server Container in the transport_url field, in the tag configuration. It instructs the Google Tag to send collected data to our Server Container, instead of the Google Analytics servers, which is the default behavior of this tag. This “Google Tag” is what we call a “Collector Tag”. In the GTM Server Container, a component that we call “Client Adaptor”, receives data, decodes it and makes it available to server tags. This Client Adaptor needs to be compatible with the Collector Tag and for this, we used the default “GA4 Client” (or simply, GA4), preinstalled on the Server Container and compatible with our Collector Tag.
Capturing traffic in the Browser. We visited our website in a new browser profile and used the Chromium debug tools to capture traffic between the Collector Tag and the Client Adaptor. We analyzed the outgoing traffic similarly to the Client-side study (§3.1.1).
Capturing Data Flows and Availability of Collected Data. To analyze the outgoing traffic of the Server Container, we instrumented our server to capture network exchanges and decrypted them by starting the Node.js interpreter, which runs the Server Container, with the --tls-keylog option (Node.js, nd). This option instructs Node.js to export encryption keys. We used Wireshark (Wireshark Foundation, nd) to collect traffic and imported the encryption keys previously exported to decrypt it. Similarly to the Client-side experiments, we identified data sent by Server Tags to Data Collectors using our server instrumentation. We finally compared our observations with the data items visible on the websites that received the user’s data.
Consent Management. To assess the ease of making the Server-side GTM architecture GDPR-compliant, we searched for official documentation on this matter and only found a relatively basic page (Google, ndg). In order to have concrete results, we added a CMP to the Web Container. We chose Cookiebot because it is compatible with Consent Mode in Client-side GTM (Server-side is not mentioned) and because it maps well to the consent variables (see Table 4). We then observed the behavior of the server tags according to the end-user consent, using the same traffic analysis methods as before.
3.1.3. Artifacts and Responsible Disclosures
Artifacts associated to this work are available in an anonymous repository (supplementary-materials, nd). Additionally, two responsible disclosures are in progress, one with Google and another one with the Consentmanager CMP. An update will be added to this article in case of feedback from the companies.
3.2. Results
In this section, we present our findings for the Client- and Server-side GTM, selected Tags, and of Google Consent Mode.
Technical finding 1 ().
Hidden data leaks by the Tags on the Client- and Server-side GTM. By analyzing data sent by the Tags and comparing with data visible to the Publisher in the Data Collector’s website, we found that certain Data Collectors do not show what data they collect. Table 1(a) shows the data sent by the three Tags that we tested in Client-side GTM without any installation of consent solutions. We observe that the data sent by the Pinterest Tag is not visible to the Publisher on the Pinterest website, where we logged in to observe Pinterest’s disclosure about collected data. Moreover, we find that the data collected by the Google Tag about form interaction is not shown in the Google Analytics dashboard. This finding demonstrates that for such Tags, Publishers are not aware of the data collected by the Tags that they select.
For Server-side GTM, Table 1(b) shows the data sent by the three selected Tags. Similarly, some collected data is not visible to the Publisher on the Data Collector’s website. The Google Analytics Tag collects data about interaction with forms and the truncated IP address of the end user. None of this is shown to the Publisher through the Google Analytics report pages. Mixpanel receives the complete IP address of the end user and URL of visited webpage, but does not show it to the Publisher. We found out that with Server-side GTM, the end-user IP address collection is necessarily deliberate: it requires the Client Adaptor to copy the IP address from the browser’s packets to data shared with all the server Tags, which is what the GA4 Client Adaptor does. In case of Mixpanel and Facebook Conversion API, moreover, the Tags send the IP address received from GA4 to Mixpanel and Facebook. Finally, Facebook did not allow us to review the collected data since when we created a new account, it was flagged as suspicious and we were blocked from accessing the data collection dashboard. Therefore, in both Client- and Server-side GTM, these hidden data leaks raise transparency concerns since Publishers are not aware of the data collection implications when they select a particular Tag. We further analyze its legal implications in Section 4.
Data Google Tag Pinterest Hotjar
URL of current page
browser and version
screen dimensions
computer architecture
operating system
OS version
engagement time
preferred language
title of current page
number of visits
forms interacted with
scrolling actions
browser windows size
clicks with position
typing in a form
submitting a form
precise mouse moves
(a) Client-side experiments.
Data Google Analytics Facebook Conversion API Mixpanel
IP address of the user
URL of current page *
browser and version
screen dimensions
computer architecture
operating system
OS version
engagement time
preferred language
title of current page
number of visits
forms interacted with
(b) Server-side experiments. Note that the Google Analytics tag truncates the last byte of the IP address claiming to anonymize it.
Table 2. Data collected by the Tags. ✓indicates the data type that is collected by the Data Collector and made visible to the Publisher via the Data Collector’s website. indicates data type that is collected but not visible, and * indicates it is sent to a blocked account.
Figure 3. Client-side GTM: The Hotjar tag downloads an arbitrary script and injects it in the page.
Technical finding 2 ().
56 Client-side Tags insert scripts that have full access to the browser APIs and page DOM. In Web Containers, tags run in a sandbox which restricts them to a limited set of the JavaScript functionalities (Google, ndk). This sandbox implements the GTM permission system that controls tags’ access to browser features, such as setting cookies or collecting other data with browser APIs. One of such permission is called “inject_script ” which allows a tag to download and execute an arbitrary script outside of the Web Container. When a tag is granted the inject_script permission, it can inject such arbitrary script, thus bypassing the GTM permission system that controls access to features, accessing to the all the browser APIs and DOM within the same origin, according to the Same-Origin Policy (mozilla.org contributors, ndc). By analyzing 78 officially supported Client-side Tags (§3.1.1), we found that 56 tags inject such scripts. Figure 3 illustrates how Hotjar Tag injects its own script in the example.com page. Within the GTM permission system, the Hotjar Tag has the permission to read and write in two global JavaScript variables (hj and hj.q) and to inject a script. We find out that it collects all the data shown in Table 1(a), including data that could not be accessed from the sandbox, such as precise mouse movements. Surprisingly, we found that 5 Tags owned by Google itself333“Google Ads”, “Google Analytics”, “Google Surveys”, “Google Tag”, “G. Trusted Stores” bypass its own GTM permission system and inject scripts in the same way as Hotjar Tag [^4] The full list of scripts is available in Supplemental Materials (supplementary-materials, nd)..
Technical finding 3 ().
GTM permission system allows injection of arbitrary scripts. By analyzing the presence of inject_script permission for all 78 Client-side Tags, we have detected 11 Tags that do not have this permission, but still inject arbitrary scripts. Out of these 11 Tags, 7 Tags are provided by Google [^5] Google owned Tags: “google tag”, “google ads calls from website conversion”, “google ads conversion tracking”, “google ads remarketing”, “google analytics: classic”, “google analytics: ga4 event”, “g. analytics: universal analytics”. Other Tags: “eulerian analytics”, “lytics js tag”, “tradedoubler lead conversion”, “tradedoubler sale conversion”.. This finding shows that the GTM permission system implemented in the Web Container sandbox allows Tags to insert arbitrary, uncontrolled scripts, thus opening potential security and privacy vulnerabilities to the website. We have disclosed this finding to Google via their Bug Bounty online system (§3.1.3).
Technical finding 4 ().
“He who says nothing agrees”: undefined consent variables are granted by default. In our experiments on consent mode and CMPs for Client-side GTM (see §3.1.1), we found that consent variables are in an undefined state when the CMP is loaded in the webpage. Though CMPs are expected to set default values to all such variables, we found out that some CMPs do not explicitly do that. Surprisingly, in this case, GTM considers all such undefined variables to be accepted by the end user, even though the end user has not interacted with the consent banner of the CMP yet. Among two CMPs tested (see §3.1.1), we detected this behavior for the Consentmanager CMP. This CMP sets a default value to only two consent variables – analytics_storage and ad_storage – leaving three GTM consent variables – security_storage , personalization_storage functionality_storage – and consent variables specific to this CMP – e.g., cmp_purpose_c56 which corresponds to the “Social Media” purpose – in undefined state. These extra variables are hence considered granted by GTM. As a result, all the Tags that depend on these four consent variables get executed even without user consent. We note that this default behavior of undefined variables cannot be changed and discuss its legal implications in the Potential violation 5 in Study II.
Technical finding 5 ().
Data collected by the Google Tag without consent. With built-in consent, tags always get executed and since they are aware of the consent choices of the user, they should adapt their behavior, and not collect any data related to purposes refused by the end user. When we tested the Google Tag, which has built-in consent with various consent decisions (accept all, decline all, only analytics, only advertisement), we found out that this Tag always sends the user’s data shown in Table 1(a), independently of the user’s consent choice. We found in the GTM documentation that when the ad_storage or analytics_storage built-in consent variables are refused, the data collected by Google is “… never used to track individual users across apps or websites, build remarketing lists, or generate user profiles” (Google, nda).
Technical finding 6 ().
Lack of consent tools in Server-side GTM. Contrary to Web Containers, the Server Container configuration interface does not provide consent configuration tools such as the association of consent variables to tags. We however found that the values of two consent variables, ad_storage and analytics_storage , are sent to our server container and provided to the Server Tags. The documentation on consent in Server Containers tells that “Google product tags in the server are consent-aware and adjust the amount and kind of data they send based on the user’s preferences.” As such, we deduce the server tags can implement a similar mechanism to built-in consent. However the Publisher cannot see on which consent variable the tag will change its behavior. Finally, for the three tags tested, we found that declining consent did not impact the transmission of data items identified in Table 1(b).
4. Study II: Legal Analysis of GTM
In this section, written together with a legal expert and co-author, we first explain the legal background on EU Data Protection law relevant to the GTM ecosystem. We then provide a legal analysis by identifying the legal role of each actor as it is crucial for further evaluation of each actor’s compliance and liability (GDPR, 2016, Recital 79).
4.1. Legal Background
The General Data Protection Regulation (GDPR) (GDPR, 2016) applies to the processing of personal data (European Data Protection Board, 2007) and imposes obligations on those actors who processes it paired with heavy fines for non-compliance. The ePrivacy Directive (ePD) (ePD-09, 2009a) provides supplementary rules to the GDPR in particular for the use of tracking technologies. Whenever cookies and other tracking technologies are stored and read from the user’s device, the ePD (ePD-09, 2009a, Art.5(3)) requires organizations to request consent for the storage of such trackers for certain purposes for processing data, such as advertising ((2013b), EDPB; (2013a), EDPB). Some purposes are exempted of consent, e.g, functional or technical purposes required for a website requested by a user to operate (ePD-09, 2009a, Recital 66). The only way to assess with certainty whether consent is required is to analyze the purpose of each tracker on a given website ((2012), EDPB; Fouad et al., 2020b).
4.1.1. Personal data
It is “any information relating to an identified or identifiable natural person (’data subject’). An identifiable natural person is one who can be identified, directly or indirectly (GDPR, 2016, Art.4(11)). In order to determine whether a person is identifiable, account should be taken of all the means likely reasonably to be used by any actor to identify that person. Accordingly, if certain data, alone, is not personal data, it becomes personal data as regards to someone who reasonably has the means of enabling that data to be associated with a specific person (Sca, 2023, para. 46). GDPR Recital 30 asserts that online identifiers provided by their devices, such as IP addresses, can be associated to a person, thus making them identifiable. This identification does not require that all the information enabling that person to be identified should be in the hands of a single entity (Bre, 2016, para. 42, 43). Processing of personal data consists of ‘any operation(s) performed on personal data, such as collecting, sharing, using, making available, accessing, combining (GDPR, 2016, Article 4(2)). In practice, this means that almost any imaginable handling of personal data constitutes processing (29 Working Party, 2010; European Court of Justice, 2023).
4.1.2. Data Controller and Data Processor
According to European Data Protection Board (EDPB) opinion (29 Working Party, 2010), to determine whether an actor is a Data Controller, its factual roles and activities have to be evaluated in a specific situation. An actor is a Data Controller if it is responsible for determining the purposes, and the means of the processing of personal data (GDPR, 2016, Art. 4(7)).
“Determines” means having the “decision-making power” (European Data Protection Board, 2020; 29 Working Party, 2010; European Court of Justice, 2018a) or “independent control” (Information Commissioner’s Office, 2018) over the purposes and means of the processing (European Data Protection Board, 2020). The Court of Justice of the EU (CJEU), Data Protection Authorities (DPAs) and the EDPB describe that such control can be derived from: professional competence (legal or implicit) (29 Working Party, 2010); factual influence, based on the real circumstances surrounding the processing; image given to data subjects and their reasonable expectations on the basis of this visibility (29 Working Party, 2010); which actororganizes, coordinates and encourages” data processing (European Court of Justice, 2018a, para 70-71); and interpretation or independent judgement exercised to perform a professional service (Information Commissioner’s Office, 2018). “Purposes” refers to “why” data is processed. Purposes need to be explicit, specified and legitimate (GDPR, 2016, Art. 5(1)(b)). “Means” refers to the how the objectives of processing are achieved. The EDPB distinguishes between “essential” and “non-essential means” and provides examples thereof  (European Data Protection Board, 2020; 29 Working Party, 2010): Essential means are closely linked to the purpose and the scope of the processing and are inherently reserved to the controller. Examples of essential means are: determining the type of personal data processed; duration of processing, recipients of personal data, or categories of data subjects. Non-essential means may be delegated to the Data Processor, involving practical implementations, like hardware or software selection, security measures, or data storage/retrieval methods.
An actor is a Data Processor when it processes personal data on behalf of the data controller (GDPR, 2016, Art. 4(8)). An actor is a processor when: i) it is dependent on the controller’s instructions regarding processing activities  (29 Working Party, 2010)(European Data Protection Board, 2020), (Art. 28(3)(a)), Recital 81); and ii) complies with those instructions (European Data Protection Board, 2020).
4.1.3. Joint Controllership
Where two or more controllers jointly determine the purposes and means of processing, they shall be joint controllers (GDPR, 2016, Art. 26(1)). Joint participation can take the form of common or converging decisions on purposes and essential means (European Data Protection Board, 2020). Decisions can be considered as converging if they complement each other and are necessary for the processing to take place in such manner that they have a tangible impact on the determination of the purposes and means of the processing. An important criterion to identify converging decisions is whether the processing would not be possible without both parties’ participation in the sense that the processing by each party is inseparable, i.e., inextricably linked (European Data Protection Board, 2020). Joint controllership requires an agreement pursuant to GDPR Article 26. However, there is no requirement for both parties to share responsibility equally (European Court of Justice, 2018b, para 43). Similarly to Data Controllers, each of the both parties do not need to have access to personal data to be considered a controller(European Court of Justice, 2018b, para 38).
4.1.4. GDPR obligations for Data controllers and Data processors
If an actor is established as a data processor, it can be held liable and fined if it fails to comply with its obligations under the GDPR (GDPR, 2016, Art. 28(3)(f), 32-36). Compliance with the GDPR is enforced by the EU Data Protection Authorities (DPAs), which monitor and supervise the application of the GDPR (GDPR, 2016, Art.55-57). Data controllers must comply with the following principles:
  • Purpose limitation: collect personal data for specific, detailed and explicit purposes and not further processed for incompatible purposes  (GDPR, 2016, (Art. 5(1)(b));
  • Minimization: collect personal data that is adequate, relevant and limited to what is necessary in relation to the purposes for which data are processed (GDPR, 2016, (Art. 5(1)(c));
  • Lawfulness: collect personal data only when it is legitimized with one legal basis (GDPR, 2016, Art. 5(1)(a)). In case this legal basis is consent, the consent must be prior to any data collection, freely given, specific, informed, unambiguous, readable, accessible, and revocable (GDPR, 2016, Art.4(11), Art. 7) (Santos et al., 2020);
  • Transparency: collect personal data when it informs the end user about the purposes, third party recipients, legal basis, among other information (GDPR, 2016, Art. 5(1)(a), 13, 14).
  • Security: ensure appropriate security of the personal data including protection against unauthorized or unlawful processing (GDPR, 2016, Art. 5(1)(f), 32).
  • Data Protection by Default: ensure that personal data is processed with the highest privacy protection so that by default personal data isn’t made accessible to an indefinite number of persons (GDPR, 2016, Art. 25(1)(2)).
  • Accountability (GDPR, 2016, Art.5(2)): the controller must be able to demonstrate compliance to each principle at any time;
4.2. Legal Analysis of Client-Side GTM
In this section we analyze the roles of the actors involved in client-side GTM: GTM Provider, CMPs, Data Collectors and Tag Providers since they raise potential violations to the GDPR and ePrivacy directive. We start our legal analysis by identifying whether these GTM actors process personal data. According to our results, both GTM Provider and Data Collector process the IP address of the end user because they receive incoming HTTP requests on their servers. Following the arguments of Santos et al. (Santos et al., 2021), if an IP address is combined with additional user data, then the IP address receiver has the means and information reasonably likely to indirectly render a user identifiable (GDPR, 2016, Recital 26). Since the Publisher’s role does not trigger potential violations, we refer to its role in the Appendix References. Figure 4 depicts the GTM actors and their relationships.
Figure 4. client-side GTM configuration by a Publisher and usage by an end user
Consent variable Description provided by GTM
ad_storage
Enables storage (such as cookies) related to advertising
analytics_storage Enables storage (such as cookies) related to analytics e.g. visit duration
functionality_storage
Enables storage that supports the functionality of the website or app e.g. language settings
personalization_storage Enables storage related to personalization e.g. video recommendations
security_storage
Enables storage related to security such as authentication functionality, fraud prevention, and other user protection
Table 3. GTM consent variables and their descriptions.
4.2.1. Consent Management Platforms (CMPs)
The website Publisher selects a CMP to manage end-user’s consent (see step (3) in Figure 4). To integrate the consent banner with GTM, the CMP needs to be compatible with the Google Consent Mode. The two CMPs tested during this study – Consentmanager and Cookiebot – offer a scanning service that automatically detects the cookies on the Publisher’s website and third parties with whom these cookies are shared (Santos et al., 2021, Fig.3). The scanner then classifies the third parties and associates purposes.
Determination of Purposes. Through this scanning activity, the CMP extracts the purposes of the trackers from the Publisher’s website as shown in step (4a) of Figure 4. From reading the provided scanning report, the Publisher configures those purposes for each tag in the GTM interface as shown in step (5a) of Figure 4. As such, the Publisher determines the purposes for processing together with the CMP’s help. Notably, CMPs are argued to be recognized as Data Controllers, following the arguments of Santos et al. (Santos et al., 2021, Section 4.3). We posit that the final decision on the needed purposes for the Tag in question is a converging decision between Publishers and chosen CMPs (see §4.1.3), where both decisions to hire the CMP by the Publisher, and to identify the purposes by the CMP complement each other and are necessary for the processing to take place. Therefore, this decision (joint decision by the Publisher and the CMP) has a tangible impact on the determination of purposes of processing.
Determination of means. When CMPs provide their services and tooling (besides consent management), such as scanning, such CMPs define the means for processing.
Legal role: Publisher and CMP are Joint Data Controllers.
Potential violation 1 ().
CMP scanners often miss purposes. During our experimentation on Client-side GTM (see §3.1.1), we installed the Google Tag that contains two specific built-in purposes – ad_storage  and analytics_storage . We found out that when Cookiebot and Consentmanager scanners run, they report only analytics purposes and miss the advertising purpose of this tag. Publishers that rely on these scanners without further reading Data Collector’s Terms&Conditions face the risk to misconfigure consent requirements of Tags. Furthermore, as third-party cookies are being phased out by all major browsers, CMP scanners only relying on cookie scanning such as Cookiebot risk under-reporting purposes for installed tags, as demonstrated by Toth et al. (Toth et al., 2022).
Recommendation CMPs should either provide a comprehensive scanning service including all purposes both from stateless and stateful tracking technologies or avoid providing any scanner at all. Relatedly, we recommend GTM to provide to the CMP/Publisher the list of tags installed in the GTM Web and Server Containers so that CMPs will be able to reliably detect the Data Collectors and map the purposes of such tags.
Cookiebot Purpose GTM consent variable(s)
Statistics analytics_storage
Marketing ad_storage
Necessary security_storage
Preferences
personalization_storage  functionality_storage
Table 4. Cookiebot purposes and corresponding GTM consent variables.
Potential violation 2 ().
Mapping CMP purposes to GTM consent variables is not compliant. CMPs need to map GTM’s consent variables to the purposes shown in their consent banner. Notably, such mappings are not documented anywhere by CMPs nor by GTM. Furthermore, these mappings are subject to the legal issues identified in Potential Violation 3 and 4. When installing CookieBot and Consentmanager CMPs (claimed to be GTM compatible (Google, ndb)) on our website, we could infer the mappings of purposes (presented in Appendix Table 4 and 8). Herein we report several problems concerning such mappings: First, these CMPs use different names for purposes rather than the ones used by GTM. For example, the consent variable analytics_storage is called “Statistics” by Cookiebot, and “Measurement” by Consentmanager. The Publisher might not be aware that the purpose “Measurement” in the Consentmanager CMP banner corresponds to the analytics_storage consent variable in the GTM interface, and might have difficulty to configure tags that depend on this purpose. Second, certain CMPs merge multiple GTM purposes altogether, as displayed in Table 4. GTM consent variables functionality_storage and personalization_storage are merged into one category called “Preferences”. A Publisher does not know that the “Preferences” purpose in the Cookiebot refers to two different purposes in GTM which are personalization_storage and functionality_storage . Third, in our experiments we ascertained that Consentmanager only partially maps its purposes to GTM consent variables, as displayed in Table 8. Consentmanager does not map any of its own purposes to the ad_storage consent variable. If a user agrees to the marketing purpose shown in the consent banner, as there is no default consent variable that is associated with marketing purpose, a new variable (not included by default in GTM) is created within GTM. A Publisher might not be aware of this lack of mapping, neither of the need to configure such consent variable to a Tag. For instance, this means that a Publisher that uses the personalization_storage variable to configure a Tag, will see that this Tag will always run independently of user choices (see Technical finding 4).
Recommendation We recommend that GTM and CMPs agree on purpose names, descriptions and matching of purposes.
4.2.2. GTM Provider
This actor [^6] The provider of GTM is Google at the time of our experiment. provides the GTM service with its Tag management functionalities to website Publishers.
Determination of purposes. This actor is involved in the determination of purposes for two reasons. First, it defines five default consent variables in GTM “Consent Mode”: ad_storage , analytics_storage , security_storage , functionality_storage , and personalization_storage , presented in Table 3. We claim that these consent variables correspond to personal data processing purposes within the meaning of Article 5(1)(b) of the GDPR. A Publisher then needs to associate each Tag to one or more of these mentioned purposes, when purposes are not built-in. Second, we argue that GTM “organizes and coordinates” data processing (European Court of Justice, 2018a, paragraphs 70, 71), since it enables the orchestration of data collection and sharing through several actors - the Publisher, Tag Provider and the Data Collector. Hence, GTM Provider’s platform is the placeholder that coordinates the Tags for collecting data. GTM also encourages data processing by proposing the tags to Publishers.
Determination of means The GTM Provider generates Web Containers and provides the configuration interface wherein the Publisher configures GTM Containers and their components (e.g., Tags). As such, it determines non-essential means for data to be processed.
Legal role: Provider of the GTM service is a Data Controller.
Potential violation 3 ().
GTM purposes are limited to client-side storage. All consent variables definitions in GTM (Table 3) are storage-based. The technologies that do not explicitly store any information in the user’s terminal equipment, known as “stateless” technologies, such as browser fingerprinting (Laperdrix et al., 2020; Fouad et al., 2022), are hence excluded and therefore not covered by the GTM Provider by default. This means that if a Publisher installs a given Tag in the GTM interface which uses a stateless technology, it will not be covered by the purposes defined by GTM. Therefore, there is no way for Publishers to be compliant with the ePrivacy Directive (ePD-09, 2009b) when using stateless tracking technologies in the Web Containers, since this directive requires consent for all non-necessary purposes of these tracking technologies. We reason that both Publishers and Tag Providers (when they decide to include built-in consent) might become confused when mapping consent variables to Tags. Because of this limitation, they cannot ensure compliance for their websites and Tags when using stateless technologies that require consent.
Recommendation GTM should define purposes regardless of the type of technology used (stateless or stateful).
Potential violation 4 ().
GTM purposes are not specific nor explicit. The GTM Provider, as a data controller, is obliged to provide specific, detailed and explicit purposes, according to the GDPR Purpose limitation principle (see Section 4.1.4). The description of the personalization_storage consent variable that says “storage related to personalization e.g. video recommendations” is not clear enough to understand to what final goal such personalization applies to - for example, if it is used for targeted advertising, it would require consent. These short and vague descriptions make it difficult for Publishers to conclude whether a purpose is exempted of consent (in case such purpose is functional or technical, see §4.1). Consequently, CMPs might run the risk to require consent for purposes exempted thereof. For example, Cookiebot requests consent for non-necessary purposes of “Preferences”, “Statistics” and “Marketing” in its consent banner. However, the “Preferences” purpose (see Table 9) seems to be necessary for a website to function, and hence, should be exempted of consent (ePD-09, 2009a, Art. 5(3))). As Cookiebot maps this purpose to the GTM consent variables personalization_storage and functionality_storage , according to their description in Table 3, it requests consent for this purpose.
Recommendation GTM should employ purpose-specific and explicit consent variables. Additionally, CMPs should not request consent for unnecessary purposes.
Potential violation 5 ().
Defaulting consent variables to “accepted” means that Tags run without consent. As described in the Technical Finding 4, the Web Container considers both “accepted” and “not defined” variables as accepted by the user, by default. We also discovered a case where the Consentmanager CMP allows some consent variables to be undefined. As such, the tags installed in the Web Container will run and collect user’s data without their due consent. This practice renders any processing of personal data potentially illegal (GDPR, 2016, Article 6(1)(a)). Notably, Consentmanager declares in its documentation that some consent mode functions are limited and not recommended for Publishers to use (Consentmanager, ndb). We argue that their disclaimer on the ”experimental mode” does not exempt them of accountability obligations (GDPR, 2016, Article 5(2)). Nevertheless, GTM still recommends to use this CMP without any warning, and it further states: “Discover featured CMP templates that deeply integrate with GTM’s consent configuration” in the interface. By doing so, GTM pushes Publishers to use a flawed CMP that might enable tags to run independently of consent, and thus enabling processing of personal data without legal basis.
Recommendation GTM should treat undefined consent variables as “not accepted” by default, and prohibit tags from collecting data before consent has been obtained to comply with the Privacy by default GDPR principle (GDPR, 2016, Art. 25(2)). GTM should validate recommended CMPs or explicitly warn about their limitations.
4.2.3. Data Collector
Data Collectors receive the data collected by tags from the Publisher’s website (Step (9a) and (9b) in Figure 4).
Collection of personal data Data Collectors have the means to combine the IP address with additional collected data, as shown in Table 2. Following our arguments in the beginning of Section 4.2, the combination of IP address with these data can render a user identifiable, and thus this data is considered to be personal data.
Determination of Purposes The Data Collector, through the Terms and Conditions (T&Cs) or Privacy Policy it provides to the Publishers, decides upon the purposes and other conditions of processing, as shown in step (4b) in Figure 4. Examples for two specific case studies on two popular Data Collectors show how they communicate purposes to Publishers through T&Cs. In the documentation for Pinterest Tag, we found only one and very specific purpose of “online behavioural advertising” in the Advertising Guidelines(Pinterest, ndb). On its page, Pinterest says: “(…) information will be shared with third parties for online behavioural advertising”. In HotJar’s documentation, we found information about the purposes of the collected data in its data processing agreement (Data Processing Agreement, nd). In our interpretation, it pertains to the purposes of analytics and aggregated user behavior: “Hotjar allows its users to analyze and understand the behavioral patterns of their visitors (…). We reason there is a pluralistic control exercised by both the Data Collector and the Publisher: there is a common and complementary decision from both Publisher and Data Collector to process data, from the very moment in which the Publisher contractualizes with the Data Collector, derives the purposes from its T&Cs (Step 4b Fig. 4) and relies on the means of processing from the Data Collector. The processing would not be possible without the participation of both parties. This interdependence of the Publisher on the definition of purposes and means of the Data Collectors renders both actors joint controllers.
Determination of Means Data Collectors receive users’ data from the Tag and provide a service to the Publisher by processing the data. For example, Google Analytics provides a dashboard with statistics about the Publisher’s website users. As a result, the Data Collector defines the means of processing since it decides how to process the received data and what algorithms to use on the received data to obtain the service requested by the Publisher.
Legal role: Publisher and Data Collector are Joint Data Controllers.
4.2.4. Tag Provider
This actor merely writes the code of a tag, and such tag consists in the means to collect and send the data to the Data Collector. In general, Tag Provider does not receive nor store data from end users. The Tag Provider is aware that, at some point, some Publisher will use its Tag, without necessarily knowing the identity of the Publisher. The Tag Provider can include built-in consent in the Tag (e.g., Google Tag in Figure 4, see “Consent Configuration” 3.1.1). As shown in step (5b) of Figure 4, GTM extracts the purposes used by every tag using built-in consent functionality and displays them in GTM consent configuration.
In Client-side GTM, it is common that a given company plays two roles at the same time: Tag Provider and Data Collector. We have identified 3 scenarios (S1, S2, S3) shown in Table 5 that we use to reason about the role of the Tag Provider.
Is a Collector?
Has built-in consent?
Legal Role of the Tag Provider
S1
none
S2 none
S3
not relevant
Data Controller
Table 5. Legal role of the Tag Provider depending on whether it is a Data Collector or the tag has built-in consent.
S1: When the company behind the Tag Provider is not the same as the Data Collector, and there is no built-in consent in the Tag , it seems that there is no contract or agreement that could link the Publisher to the Tag Provider. The Tag Provider still defines, in the Tag, to whom the user’s data will be sent (so, who are the Data Collectors). This entails that the Tag Provider determines essential means of processing. However, the Tag Provider is not a Data Controller because it is missing the determination of purposes, i.e., there is no built-in consent in the Tag that pre-defines purposes. The Tag Provider is neither a Data Processor, since it does not follow instructions from the Publisher nor from the Data Collector – it does not process personal data on behalf of these actors, since it is missing the dependence on the controller’s instructions regarding processing activities which defines the role between a processor and a controller (see §4.1). Hence, Tag Providers, under this scenario, are just service providers, with no GDPR obligations.
S2: When the company behind the Tag Provider is not the same as the Data Collector, and the tag contains built-in consent, the Tag Provider includes the purposes for collecting data inside of the Tag’s code. This inclusion of predefined purposes in the Tag could qualify the Tag Provider as a Data Controller, since it seems to define the purposes itself. Even though the Tag Provider develops a data-collecting Tag by itself, which is released to the public and enables the collection of data, these facts are not enough to reveal a factual influence over the definition of purposes of personal data (AG-, 2023). We reason that by embedding the built-in consent in the tag, the Tag Provider might not relate to the processing of personal data in itself; it just consists of a prior step to the processing of personal data (AG-, 2023), it does not provide enough elements of a factual nature that the Tag Provider exercises an actual influence with regards to the “purposes and means” of processing data (AG-, 2023). Also, this actor does not process personal data. Hence, Tag Providers, under this scenario, are just service providers, with no GDPR obligations.
S3: When the company behind a given Tag Provider is the same as a Data Collector it is possible to conclude that a Tag Provider/Data Collector, jointly with a given Publisher, can be considered as Joint Controllers, since Data Collectors and Publishers are Joint Controllers, as already established in Section A.1. This qualification holds regardless of whether there is built-in consent or not in the Tag, since the Data Collector determines already the purposes of processing in the T&C or contractual services.
Legal role: Tag Provider’s role depends on whether it is a Data Collector. See Table 5 for a summary of their role.
Potential violation 6 ().
Google Tag sends data independently of user’s consent decisions. Our Technical Finding 5, Google Tag always sends user’s data to Google even when the user has rejected all built-in purposes, thereby processing user’s data without a lawful legal basis, infringing the Lawfulness principle. Moreover, when user’s data is sent to Google, the Security principle is violated because of an unauthorized disclosure with Google (§4.1). As a Data Controller, Google is liable for these potential violations.
Recommendations Google, as Tag Provider and Data Collector, must change its “Google Tag” behavior to respect user’s choices in order to be compliant.
Potential violation 7 ().
GTM allows Tag Providers to inject scripts exposing end users to security risks. In Technical Finding 2, we found that GTM contains a special inject_script permission, allowing Tags to inject arbitrary JavaScript code in the website’s page (see Figure 3). We found 56 Client-side Tags that inject such scripts which are not subject to any security measures implemented in Web browsers, such Same-Origin Policy (mozilla.org contributors, ndc), to protect users from script-based Web attacks. This practice, allowed by GTM and executed by Tags, potentially infringes the Security principle (§4.1.4). The Provider of the GTM service, as a Data Controller, is liable for these potential violations. Additionally, the Tag Providers of these 56 Tags are Data Collectors, when they are Data Controllers, are also liable for such violations.
Recommendations GTM should not allow Tag Providers to include arbitrary scripts in the website’s pages without adopting any security safeguards. Tag Providers should not use this loophole of GTM to inject arbitrary scripts.
4.3. Legal analysis of Server-Side GTM
In this section we describe the actors involved in Server-side GTM. We reproduce the same legal reasoning about some of the roles of the actors held on Section 4.2 regarding the Publisher (appendix §A.1), Data Collector (§4.2.3) and GTM Service Provider (§4.2.2). As such, herein we focus on the Provider of Collector Tag and Client Adaptor, the Server Provider and Server Tag Provider. Since the Server Provider’s role does not trigger potential violations, we refer to it in the Appendix (A.2). After analyzing the Provider of Collector Tag and Client Adaptor functions, we concluded it has no GDPR legal role. We found that this actor presents similarities with the Tag Provider in client-side GTM (when it is not a Data Collector) since it merely develops and provides software without processing personal data and without following instructions from any other actors. Table 6 outlines server-side roles for each GTM actor.
The processing of personal data on the Server-side needs to be assessed on a case-by-case basis. The Server Provider processes the IP address of end users since it receives incoming HTTP requests on its servers; the Data Collector will sometimes receive personal information (see Table 1(b) for specific Data Collectors that receive the IP address of end users).
4.3.1. Server Tag Provider
This actor presents similar functions as to the Tag Provider on client side 4.2.4. On server-side GTM, the absence of the built-in function does not change its legal role as a Data Controller when the Server Tag Provider is also a Data Collector.
Legal role: Tag Provider’s legal role depends on whether the Data Collector is the same company.
Potential violation 8 ().
Server Tag Providers that are also Data Collectors are aware that lawful data collection is not possible. A Server Tag Provider, who provides the code of a tag, knows how the GTM Server Container functions (see Figure 1(b)) and thus is aware that no consent management tool is available on the Server Container. A Server Tag Provider also knows that the Publisher cannot conveniently configure the Server Container to match GTM consent variables to Server Tags (see §5). Nonetheless, Server Tag Provider still provides Server Tags. When a Server Tag Provider directly receives users’ personal data, it has a legal role of a Data Collector, and therefore becomes a Data Controller (see Table 5). Moreover, it is a Joint Controller with the Publisher (see §4.2.3), and therefore has the responsibility to ensure lawful data processing.
Recommendation Server Tag Providers that are Data Collectors should not provide Server Tags until the GTM does not have the consent management functionality on the Server-side.
Actor Client-side Server-side
Publisher Data Controller Data Controller
Data Collector
Data Controller Data Controller
CMP Data Controller N/A
Tag Provider
Depends, see Tab.5 Depends, see Tab.5
GTM Provider Data Controller Data Controller
Server Provider
N/A Data Processor
Provider of Collector Tag and Client Adaptor N/A No legal role
Table 6. Summary of the legal roles of the actors.
5. Discussion
In this section, we discuss other findings and reflect upon the difficulty to comply with the EU Data Protection framework for various actors within the Client- and Server-side GTM and provide further recommendations for improvement of GTM.
Complying with data subject rights is hard for the Publisher
Within both Client- and Server-Side GTM, the Publisher is left alone to comply with users’ rights, like data requests, due to the absence of a dedicated system for this functionality. For example, if a user requests access to her data (under Article 15 GDPR), the Publisher would probably need to find the contact of every Data Collector and compile it manually to then answer to the user – a task confirmed to be problematic as studied by Samarin et al. (Samarin et al., 2023). GTM should furnish Publishers with a common interface to identify all Data Collectors, and a streamlined tool from Data Collectors to facilitate users’ data requests.
Built-in consent raises trust issues
When the Publisher uses Tags with built-in consent (such as “Google Tag” in Potential Violation 6), it relies on the Tag Provider to properly implement the built-in consent in its code. However, the Tag Provider can declare to rely on some built-in consent purposes, but completely ignore them in the Tag code, and send the user’s data even if the corresponding purposes were rejected by the user. The only way a Publisher can ensure proper consent management is to review the Tag’s code. This in only possible for non-official Tags, whose code is available on the template gallery. However, code reviewing is not possible for official Tags, that are implemented as sandboxed code inside the gtm.js , and it requires heavy reverse engineering. Publishers can also ignore built-in consent and manually add consent variables (Step (5a) in Figure 4). In this case, the Tag will only run when all the consent variables associated with the tag are granted. GTM should allow Publishers to access the code of the Tags for auditing and should facilitate Publishers to understand the Tags’ code.
Server-side GTM is invisible for regulatory monitoring and auditing.
Server-side GTM obstructs compliance auditing endeavors from regulators, data protection officers, and researchers, since data collection occurs remotely on a server, whereas traditionally it happens in the browser. For example, while only the Collector Tag is detectable, the tags chosen by a Publisher remain invisible. It is possible to monitor the data collected and transmitted to the GTM Server Container; however, discerning which Collectors have access to such data remains unknown. Moreover, auditing and monitoring is exclusively attainable by only contacting the Publisher to grant access to the configuration of the GTM Server Container. Furthermore, the Publisher is able to change the configuration of the GTM Server Container at any point in time (e.g., before any regulatory investigation), masking any compliance check.
Consent is hard to configure on GTM Server Containers.
While Google advertises: “respect user consent choices with Google Tag Manager” (Google, ndj), consent management tools are limited to client-side GTM and it is challenging to render consent compliant on the server-side. The lack of consent management tools in GTM Server Container raises the following two problems: First, CMPs scanners can only detect tags in GTM Web Containers but are not able to detect tags on GTM Server Containers; this impossibility entails that CMPs cannot display the purposes and the Data Collectors in the consent banner to end users. This type of information disclosure is mandatory for an end-user informed consent and transparent processing of personal data (GDPR, 2016, Art. 4(11), 5(1)(a)). Second, Publishers that wish to be compliant on the Server-side, need to make complex configurations such as writing code to decode consent information and create filtering rules in respect to the given consent in the GTM Server Container since there is no interface to configure consent. Furthermore, since Server Container does not support consent mode, Publishers cannot associate tags with purposes. GTM should implement consent management tools in GTM Server Containers and provide the list of installed Tags to CMPs.
6. Conclusion
This work is the first to study both versions of Google Tag Manager: Client-side and Server-side. We analyzed these systems with 78 Client-side Tags, 8 Server-side Tags and two Consent Management Platforms (CMPs) and performed an in-depth technical and legal analysis of GTM, determining the responsibilities and potential legal violations of each actor. Our results show that GTM has many pitfalls, such as flaws in its security system and non-compliant defaults. We conclude that GTM in its current state introduces more legal issues than solving, while making compliance difficult to achieve for various actors and complex for regulators to monitor.
Acknowledgements.
This work has been supported by the ANR 22-PECY-0002 IPoP (Interdisciplinary Project on Privacy) project of the Cybersecurity PEPR. The authors would like to thank Javiera Bermudez Alegria, from Universidad de Chile, who contributed to the initial work.
Appendix A Actors
A.1. Publisher on the Client and Server-side
The Publisher uses GTM to include a dedicated service in their website. When a Publisher chooses the Data Collector or its tag as shown in step (1) of Figure 4, it needs to accept the Terms and Conditions (T&Cs) related to this contractual service of the Data Collector.
Determination of purposes
The selection of the Data Collector has a determinant consequence on the qualification of the Publisher in the determination of purposes. Firstly, the Data Collector establishes the purposes (Step (4b) in Figure 4) for processing data because it explicitly mentions them (e.g. analytics, among others) in the T&C or other policy documents. Since the Publisher agrees to the purposes defined in the T&Cs by the Data Collector, these are the purposes that the Publisher presents to the end-users. This presentation of purposes given to data subjects, and their reasonable expectations on the basis of this visibility (29 Working Party, 2010), reflects the determinant power that Publishers have regarding purposes.
Determination of means
Because of the acceptance of the Data Collector’s T&C, the Publisher agrees with the means of the processing that were established by the Collector, such as how to process the received data, what algorithms to use and what code to run on the received data. Accordingly, the Publisher determines the means for data processing.
Legal role: Publisher is a Controller.
A.2. Server Provider on Server-side
This actor provides to the Publisher the hardware for a given software (e.g. the GTM Server Container) to run. Two types of Server Providers exist: i) the “GTM-server service provider” which pre-installs the GTM server software, and also manages the security of the server; and ii) the “simple server provider” that provides only a generic server, and allows any given Publisher to install the GTM server software themselves. In this paper we only analysed point ii), a “simple server provider”. The service provided (i.e. the hardware for the server to run) is usually a service a Publisher needs to pay for. The Publisher hires a given Server Provider to provide the server and a contract is performed between the Publisher and such Server Provider. The Publisher will use the server to install the GTM Server Container software and consequently gives instructions the this Server on how to run it in order to process personal data. Accordingly, there is a functional dependence of the Server Provider on the Publisher’s instructions with renders this actor a data processor.
Determination of purposes
By only providing the infrastructural hardware service, the Server Provider does not determine purposes.
Determination of means
The Server Provider provides the means for a Publisher to run the GTM Server Container software, such as the hardware used and security of the system. These factual activities regarding the practical implementation of the server consist of “non-essential means”. Since the hardware provided by the Server Provider runs the GTM Server Container software – which processes personal data –the Server Provider is a data processor.
Legal role: Server Provider is a Data Processor
We did not find compliance issues of this actor.
Appendix B Client side Terminology
A Tag is a JavaScript library usually developed by marketing companies, analytics companies, independent developers or, less frequently, publishers themselves. The role of a tag is to collect information about the user or its device and send the collected data to a third-party service. In GTM specifically, a tag is a software component written in “sandboxed JavaScript” which is a limited version of JavaScript created by Google (Google, ndk). In GTM, Tags also provide functionalities (e.g., implement consent banners). GTM has a permission system for tags (Google, nde) to let them access to functionalities such as sending requests to third party servers or access cookies.
A GTM Web Container is a group of tags and their associated configuration rules. Technically, a container is a JavaScript program generated automatically by Google and downloaded by the user’s browser as a file called gtm.js from the https://googletagmanager.com. It embeds the tags selected by the publisher and their configuration as well as an interpreter that executes the “sandboxed JavaScript” code the tags are written in.
Container configuration is the process by which the Publisher selects the tags and configures them.using the configuration web interface, provided by Google at https://tagmanager.google.com.
Consent Variables store the consent of the end user in the GTM Web Container. They can be in three states: undefined, granted or denied. In examples given by GTM in the documentation, consent variables are used to represent purposes for data collection and processing. For instance, a hypothetical “targeted_advertising” consent variable could represent the consent or refusal of the end user to let their data be used to display targeted ads. Tags can be configured to have different behaviors (regarding data collection) depending on the state of these variables.
Built-in consent is a feature that tags can implement. A tag supports this feature is able to adjust its behavior based on the state of consent variables. For instance, when a end user only accepts that their data to be used for analytics purposes but not for advertisement purposes, a tag supporting built in consent and collecting data for both purposes can limit its data collection to analytic purposes only. Technically, a tag implementing built-in consent uses the GTM provided isConsentGranted API to check the status (accepted or declined) of each of the purposes.We provide an example of code of a tag implementing built-in consent in Appendix Figure LABEL:fig:builtin-consent-code-example.
Manual consent configuration is the process by which the Publisher associate consent variables to the tags installed in the GTM Web Container. A tag that is manually associated with one or more consent variables will be prevented from running by the GTM Web Container until all its associated variables are set to granted. For instance, in this case, when a user only accepts that their data to be used for analytics purposes but not for advertisement purposes, a tag configured with manual consent for both analytics and advertisement purposes will not run.
Appendix C Server side Terminology
A Server Tag is a component written in “sandboxed JavaScript” (Google, ndk) that runs exclusively on the server. It can process and send the data that is received, by the GTM Server Container, to a third party server. Server Tags can store first-party cookies in the user’s browser.
A Client adaptor is a plugin, embedded in the server container. It decodes data arriving to the GTM server Container and converts them into events.
An Event is something that happened (e.g., the user clicked on a button) and it’s associated data (e.g., the button’s name, its position on the page).
A Collector tag tag is a tag running in the browser that collects data and send it through the network to a GTM Server Container.
Cookie detected Purpose associated Tag creating the cookie
is_eu
Necessary
Pinterest
CookieConsent
Necessary
Cookiebot
_hjRecordingEnabled
Statistics
Hotjar
_hjRecordingLastActivity
Statistics
Hotjar
hjViewportId
Statistics
Hotjar
hjActiveViewportIds
Statistics
Hotjar
_hjCookieTest
Statistics
Hotjar
_ga_#
Statistics
Google Analytics
_ga
Statistics
Google Analytics
_hjSessionUser_#
Statistics
Hotjar
_hjFirstSeen
Statistics
Hotjar
_hjIncludedInSessionSample_#
Statistics
Hotjar
_hjSession_#
Statistics
Hotjar
_hjAbsoluteSessionInProgress
Statistics
Hotjar
ar_debug
Marketing
Pinterest
_pinterest_ct_ua
Marketing
Pinterest
v3/
Marketing
Pinterest
_pin_unauth
Marketing
Pinterest
Table 7. Results of the Cookiebot scanner. We deduct which tag that is responsible for the cookie creation by reading the description of the cookie and the provider shown in the report.
Consentmanager purpose GTM consent variable
Measurement analytics_storage
N/A ad_storage
N/A security_storage
N/A personalization_storage
N/A functionality_storage
Marketing N/A
Function N/A
Preferences N/A
Other N/A
Social Media N/A
Table 8. Correspondence between Consentmanager purposes and GTM purposes. Not all GTM purposes are used by Consentmanager and some Consentmanager purposes are not mapped to GTM consent mode purposes.
Cookiebot Purpose Description
Necessary
Necessary cookies help make a website usable by enabling basic functions like page navigation and access to secure areas of the website. The website cannot function properly without these cookies.
Preferences
Preference cookies enable a website to remember information that changes the way the website behaves or looks, like your preferred language or the region that you are in.
Statistics
Statistic cookies help website owners to understand how visitors interact with websites by collecting and reporting information anonymously.
Marketing
Marketing cookies are used to track visitors across websites. The intention is to display ads that are relevant and engaging for the individual user and thereby more valuable for publishers and third party advertisers.
Table 9. Purposes defined by Cookiebot with their description
Figure 5. Summary of the legal role of the Publisher as a Joint Controller with other actors of the GTM ecosystem. In Figure 5 we summarise the legal roles of the Publisher in relation to: i) the Collector (always join controllers), ii) CMPs (joint controllers), iii) Tag Providers (when the Tag Provider is the Data Collector).
Figure 6. Confirmation window when installing a tag from the Community template Gallery. Permissions (or capabilities) of the tag are explicitely listed.

### 22 Dec 2023 v2

@@
 Artifacts associated to this work are available in an anonymous repository (supplementary-materials, nd).
-Additionally, two responsible disclosures are in progress, one with Google and another one with the Consentmanager CMP.
-An update will be added to this article in case of feedback from the companies.
+Additionally, one responsible disclosure is in progress with Google.
+An update will be added to this article in case of feedback from the company.
 3.2. Results
@@
 “He who says nothing agrees”: undefined consent variables are granted by default.
-In our experiments on consent mode and CMPs for Client-side GTM (see §3.1.1), we found that consent variables are in an undefined state when the CMP is loaded in the webpage.
-Though CMPs are expected to set default values to all such variables, we found out that some CMPs do not explicitly do that.
+In our experiments on consent mode and CMPs for Client-side GTM (see §3.1.1), we found that consent variables are in an undefined state before the CMP is loaded in the webpage.
+Though CMPs are expected to set a default value to all such variables before the tags load, we found out that some CMPs do not do it fast enough.
 Surprisingly, in this case, GTM considers all such undefined variables to be accepted by the end user, even though the end user has not interacted with the consent banner of the CMP yet.
-Among two CMPs tested (see §3.1.1), we detected this behavior for the Consentmanager CMP.
-This CMP sets a default value to only two consent variables – analytics_storage and ad_storage – leaving three GTM consent variables – security_storage , personalization_storage functionality_storage – and consent variables specific to this CMP – e.g., cmp_purpose_c56 which corresponds to the “Social Media” purpose – in undefined state.
-These extra variables are hence considered granted by GTM.
-As a result, all the Tags that depend on these four consent variables get executed even without user consent.
-We note that this default behavior of undefined variables cannot be changed and discuss its legal implications in the Potential violation 5 in Study II.
+Among two CMPs tested (see §3.1.1), we detected this behavior for the Consentmanager CMP, which correctly sets the analytics_storage and ad_storage before tag loading but define other CMP-specific variables such as cmp_purpose_c56 – which corresponds to the “Social Media” purpose – after tags already loaded.
+Since such variables are undefined when the tags are run, GTM considers them granted.
+As a result, all the Tags that depend on these variables get executed even without user consent.
+We note that this default behavior of undefined variables cannot be changed by the Publisher and discuss its legal implications in the Potential violation 5 in Study II.
 Technical finding 5 ().
@@
 The ePrivacy Directive (ePD) (ePD-09, 2009a) provides supplementary rules to the GDPR in particular for the use of tracking technologies.
-Whenever cookies and other tracking technologies are stored and read from the user’s device, the ePD (ePD-09, 2009a, Art.5(3)) requires organizations to request consent for the storage of such trackers for certain purposes for processing data, such as advertising ((2013b), EDPB; (2013a), EDPB).
+Whenever cookies and other tracking technologies are stored and read from the user’s device, the ePD (ePD-09, 2009a, Art.
+5(3)) requires organizations to request consent for the storage of such trackers for certain purposes for processing data, such as advertising ((2013b), EDPB; (2013a), EDPB).
 Some purposes are exempted of consent, e.g, functional or technical purposes required for a website requested by a user to operate (ePD-09, 2009a, Recital 66).
@@
 It is “any information relating to an identified or identifiable natural person (’data subject’).
-An identifiable natural person is one who can be identified, directly or indirectly (GDPR, 2016, Art.4(11)).
+An identifiable natural person is one who can be identified, directly or indirectly (GDPR, 2016, Art.
+4(11)).
 In order to determine whether a person is identifiable, account should be taken of all the means likely reasonably to be used by any actor to identify that person.
@@
 42, 43).
-Processing of personal data consists of ‘any operation(s) performed on personal data, such as collecting, sharing, using, making available, accessing, combining (GDPR, 2016, Article 4(2)).
+Processing of personal data consists of ‘any operation(s) performed on personal data, such as collecting, sharing, using, making available, accessing, combining (GDPR, 2016, Art.
+4(2)).
 In practice, this means that almost any imaginable handling of personal data constitutes processing (29 Working Party, 2010; European Court of Justice, 2023).
 4.1.2. Data Controller and Data Processor
-According to European Data Protection Board (EDPB) opinion (29 Working Party, 2010), to determine whether an actor is a Data Controller, its factual roles and activities have to be evaluated in a specific situation.
+According to European Data Protection Board (EDPB) opinion (29 Working Party, 2010, para 20), to determine whether an actor is a Data Controller, its factual roles and activities have to be evaluated in a specific situation.
 An actor is a Data Controller if it is responsible for determining the purposes, and the means of the processing of personal data (GDPR, 2016, Art.
@@
 “Determines” means having the “decision-making power” (European Data Protection Board, 2020; 29 Working Party, 2010; European Court of Justice, 2018a) or “independent control” (Information Commissioner’s Office, 2018) over the purposes and means of the processing (European Data Protection Board, 2020).
-The Court of Justice of the EU (CJEU), Data Protection Authorities (DPAs) and the EDPB describe that such control can be derived from: professional competence (legal or implicit) (29 Working Party, 2010); factual influence, based on the real circumstances surrounding the processing; image given to data subjects and their reasonable expectations on the basis of this visibility (29 Working Party, 2010); which actororganizes, coordinates and encourages” data processing (European Court of Justice, 2018a, para 70-71); and interpretation or independent judgement exercised to perform a professional service (Information Commissioner’s Office, 2018).
+The Court of Justice of the EU (CJEU), Data Protection Authorities (DPAs) and the EDPB describe that such control can be derived from: professional competence (legal or implicit) (29 Working Party, 2010); factual influence, based on the real circumstances surrounding the processing; image given to data subjects and their reasonable expectations on the basis of this visibility (29 Working Party, 2010); which actororganizes, coordinates and encourages” data processing (European Court of Justice, 2018a, para.
+70-71); and interpretation or independent judgement exercised to perform a professional service (Information Commissioner’s Office, 2018).
 “Purposes” refers to “why” data is processed.
@@
 4(8)).
-An actor is a processor when: i) it is dependent on the controller’s instructions regarding processing activities  (29 Working Party, 2010)(European Data Protection Board, 2020), (Art.
-28(3)(a)), Recital 81); and ii) complies with those instructions (European Data Protection Board, 2020).
+An actor is a processor when: i) it is dependent on the controller’s instructions regarding processing activities  (29 Working Party, 2010)(European Data Protection Board, 2020),  (GDPR, 2016, Art.
+28(3)(a), Recital 81 )); and ii) complies with those instructions (European Data Protection Board, 2020).
 4.1.3. Joint Controllership
@@
 26(1)).
-Joint participation can take the form of common or converging decisions on purposes and essential means (European Data Protection Board, 2020).
+Joint participation can take the form of common or converging decisions on purposes and essential means (European Data Protection Board, 2020) (paragraph 51).
 Decisions can be considered as converging if they complement each other and are necessary for the processing to take place in such manner that they have a tangible impact on the determination of the purposes and means of the processing.
 An important criterion to identify converging decisions is whether the processing would not be possible without both parties’ participation in the sense that the processing by each party is inseparable, i.e., inextricably linked (European Data Protection Board, 2020).
-Joint controllership requires an agreement pursuant to GDPR Article 26.
-However, there is no requirement for both parties to share responsibility equally (European Court of Justice, 2018b, para 43).
-Similarly to Data Controllers, each of the both parties do not need to have access to personal data to be considered a controller(European Court of Justice, 2018b, para 38).
+Joint controllership requires an agreement pursuant to GDPR Article 26 (GDPR, 2016).
+However, there is no requirement for both parties to share responsibility equally (European Court of Justice, 2018b, para.
+43).
+Similarly to Data Controllers, each of the both parties do not need to have access to personal data to be considered a controller(European Court of Justice, 2018b, para.
+38).
 4.1.4. GDPR obligations for Data controllers and Data processors
@@
 28(3)(f), 32-36).
-Compliance with the GDPR is enforced by the EU Data Protection Authorities (DPAs), which monitor and supervise the application of the GDPR (GDPR, 2016, Art.55-57).
+Compliance with the GDPR is enforced by the EU Data Protection Authorities (DPAs), which monitor and supervise the application of the GDPR (GDPR, 2016, Art.
+55-57).
 Data controllers must comply with the following principles:
@@
 5(1)(a)).
-In case this legal basis is consent, the consent must be prior to any data collection, freely given, specific, informed, unambiguous, readable, accessible, and revocable (GDPR, 2016, Art.4(11), Art.
-7) (Santos et al., 2020);
-  • Transparency: collect personal data when it informs the end user about the purposes, third party recipients, legal basis, among other information (GDPR, 2016, Art. 5(1)(a), 13, 14).
-  • Security: ensure appropriate security of the personal data including protection against unauthorized or unlawful processing (GDPR, 2016, Art. 5(1)(f), 32).
+In case this legal basis is consent, the consent must be prior to any data collection, freely given, specific, informed, unambiguous, readable, accessible, and revocable (GDPR, 2016, Art.
+4(11), 7) (Santos et al., 2020);
+  • Transparency: collect personal data when it informs the end user about the purposes, third party recipients, legal basis, among other information (GDPR, 2016, Art. 5(1)(a), 13-14).
+  • Security: ensure appropriate security of the personal data including protection against unauthorized or unlawful processing (GDPR, 2016, Art. 5(1)(f), 32);
   • Data Protection by Default: ensure that personal data is processed with the highest privacy protection so that by default personal data isn’t made accessible to an indefinite number of persons (GDPR, 2016, Art.
-25(1)(2)).
-  • Accountability (GDPR, 2016, Art.5(2)): the controller must be able to demonstrate compliance to each principle at any time;
+25(1)(2));
+  • Accountability: the controller must be able to demonstrate compliance to each principle at any time (GDPR, 2016, Art. 5(2)).
 4.2. Legal Analysis of Client-Side GTM
@@
 To integrate the consent banner with GTM, the CMP needs to be compatible with the Google Consent Mode.
-The two CMPs tested during this study – Consentmanager and Cookiebot – offer a scanning service that automatically detects the cookies on the Publisher’s website and third parties with whom these cookies are shared (Santos et al., 2021, Fig.3).
+The two CMPs tested during this study – Consentmanager and Cookiebot – offer a scanning service that automatically detects the cookies on the Publisher’s website and third parties with whom these cookies are shared (Santos et al., 2021, Fig.
+3).
 The scanner then classifies the third parties and associates purposes.
@@
 Notably, CMPs are argued to be recognized as Data Controllers, following the arguments of Santos et al.
-(Santos et al., 2021, Section 4.3).
+(Santos et al., 2021, §4.3).
 We posit that the final decision on the needed purposes for the Tag in question is a converging decision between Publishers and chosen CMPs (see §4.1.3), where both decisions to hire the CMP by the Publisher, and to identify the purposes by the CMP complement each other and are necessary for the processing to take place.
@@
 (Toth et al., 2022).
-Recommendation CMPs should either provide a comprehensive scanning service including all purposes both from stateless and stateful tracking technologies or avoid providing any scanner at all.
+Recommendation CMPs should provide a comprehensive scanning service including all purposes both from stateless and stateful tracking technologies.
 Relatedly, we recommend GTM to provide to the CMP/Publisher the list of tags installed in the GTM Web and Server Containers so that CMPs will be able to reliably detect the Data Collectors and map the purposes of such tags.
@@
 Potential violation 2 ().
-Mapping CMP purposes to GTM consent variables is not compliant.
-CMPs need to map GTM’s consent variables to the purposes shown in their consent banner.
-Notably, such mappings are not documented anywhere by CMPs nor by GTM.
+Mapping between CMP purposes to GTM consent variables is not compliant.
+CMPs map the purposes shown in the CMP banner to consent variables in GTM.
+Notably, such mappings are not documented anywhereby Cookiebot nor GTM.
 Furthermore, these mappings are subject to the legal issues identified in Potential Violation 3 and 4.
-When installing CookieBot and Consentmanager CMPs (claimed to be GTM compatible (Google, ndb)) on our website, we could infer the mappings of purposes (presented in Appendix Table 4 and 8).
+When installing CookieBot (claimed to be GTM compatible (Google, ndb)) on our website, we could infer the mappings of purposes 8).
 Herein we report several problems concerning such mappings: First, these CMPs use different names for purposes rather than the ones used by GTM.
@@
 First, it defines five default consent variables in GTM “Consent Mode”: ad_storage , analytics_storage , security_storage , functionality_storage , and personalization_storage , presented in Table 3.
-We claim that these consent variables correspond to personal data processing purposes within the meaning of Article 5(1)(b) of the GDPR.
+We claim that these consent variables correspond to personal data processing purposes within the meaning of Article 5(1)(b) of the GDPR (GDPR, 2016).
 A Publisher then needs to associate each Tag to one or more of these mentioned purposes, when purposes are not built-in.
-Second, we argue that GTM “organizes and coordinates” data processing (European Court of Justice, 2018a, paragraphs 70, 71), since it enables the orchestration of data collection and sharing through several actors - the Publisher, Tag Provider and the Data Collector.
+Second, we argue that GTM “organizes and coordinates” data processing (European Court of Justice, 2018a, para.
+70-71), since it enables the orchestration of data collection and sharing through several actors - the Publisher, Tag Provider and the Data Collector.
 Hence, GTM Provider’s platform is the placeholder that coordinates the Tags for collecting data.
@@
 GTM purposes are not specific nor explicit.
-The GTM Provider, as a data controller, is obliged to provide specific, detailed and explicit purposes, according to the GDPR Purpose limitation principle (see Section 4.1.4).
+The GTM Provider, as a data controller, is obliged to provide specific, detailed and explicit purposes, according to the GDPR Purpose limitation principle (see §4.1.4).
 The description of the personalization_storage consent variable that says “storage related to personalization e.g.
@@
 As Cookiebot maps this purpose to the GTM consent variables personalization_storage and functionality_storage , according to their description in Table 3, it requests consent for this purpose.
-Recommendation GTM should employ purpose-specific and explicit consent variables. Additionally, CMPs should not request consent for unnecessary purposes.
+Among the two CMPs we have tested, this risk only affects the Cookiebot CMP.
+Recommendation GTM should employ purpose-specific and explicit consent variables. Additionally, Cookiebot should not request consent for unnecessary purposes.
 Potential violation 5 ().
@@
 As such, the tags installed in the Web Container will run and collect user’s data without their due consent.
-This practice renders any processing of personal data potentially illegal (GDPR, 2016, Article 6(1)(a)).
+This practice renders any processing of personal data potentially illegal (GDPR, 2016, Art.
+6(1)(a)).
 Notably, Consentmanager declares in its documentation that some consent mode functions are limited and not recommended for Publishers to use (Consentmanager, ndb).
-We argue that their disclaimer on the ”experimental mode” does not exempt them of accountability obligations (GDPR, 2016, Article 5(2)).
+We argue that such disclaimer does not exempt CMPs of accountability obligations (GDPR, 2016, Art.
+5(2)).
 Nevertheless, GTM still recommends to use this CMP without any warning, and it further states: “Discover featured CMP templates that deeply integrate with GTM’s consent configuration” in the interface.
-By doing so, GTM pushes Publishers to use a flawed CMP that might enable tags to run independently of consent, and thus enabling processing of personal data without legal basis.
+By doing so, GTM pushes Publishers to use a CMP that might enable tags to run independently of consent, and thus enabling processing of personal data without legal basis.
 Recommendation GTM should treat undefined consent variables as “not accepted” by default, and prohibit tags from collecting data before consent has been obtained to comply with the Privacy by default GDPR principle (GDPR, 2016, Art.
@@
 In HotJar’s documentation, we found information about the purposes of the collected data in its data processing agreement (Data Processing Agreement, nd).
-In our interpretation, it pertains to the purposes of analytics and aggregated user behavior: “Hotjar allows its users to analyze and understand the behavioral patterns of their visitors (…).
+In our interpretation, it pertains to the purposes of analytics and aggregated user behavior: “Hotjar allows its users to analyze and understand the behavioral patterns of their visitors (…)”.
 We reason there is a pluralistic control exercised by both the Data Collector and the Publisher: there is a common and complementary decision from both Publisher and Data Collector to process data, from the very moment in which the Publisher contractualizes with the Data Collector, derives the purposes from its T&Cs (Step 4b Fig.
@@
 This inclusion of predefined purposes in the Tag could qualify the Tag Provider as a Data Controller, since it seems to define the purposes itself.
-Even though the Tag Provider develops a data-collecting Tag by itself, which is released to the public and enables the collection of data, these facts are not enough to reveal a factual influence over the definition of purposes of personal data (AG-, 2023).
+Even though the Tag Provider develops a data-collecting Tag by itself, which is released to the public and enables the collection of data, these facts are not enough to reveal a factual influence over the definition of purposes of personal data (AG-, 2023, par.
+32).
 We reason that by embedding the built-in consent in the tag, the Tag Provider might not relate to the processing of personal data in itself; it just consists of a prior step to the processing of personal data (AG-, 2023), it does not provide enough elements of a factual nature that the Tag Provider exercises an actual influence with regards to the “purposes and means” of processing data (AG-, 2023).
@@
 4.3.1. Server Tag Provider
-This actor presents similar functions as to the Tag Provider on client side 4.2.4.
+This actor presents similar functions as to the Tag Provider on client side (see §4.2.4).
 On server-side GTM, the absence of the built-in function does not change its legal role as a Data Controller when the Server Tag Provider is also a Data Collector.